OpenStudy (anonymous):

find the degree and radian measure of the angle in standard position formed by rotating the terminal side of 1/9 of a circle. How do I set up this up?

OpenStudy (ranga):

Full circle is 360 degrees. Therefore, 1/9 of a circle = ? (how many degrees?)

OpenStudy (anonymous):

Thats what their asking

OpenStudy (ranga):

Well, if a full circle forms 360 degrees then one-ninth of a circle will form one-ninth of 360 degrees. Can you tell me what is one-ninth of 360 degrees?

OpenStudy (ranga):

1/9 x 360 = ?

OpenStudy (ranga):

In degrees: 360 / 9 = 40 degrees In radians: A full circle has 2(pi) radians. Therefore, one-ninth of a circle will have 2/9(pi) radians. 

OpenStudy (anonymous):

ohh ok thanks

OpenStudy (ranga):

np